OpenStudy (anonymous):

A clerk takes h to process 9 applications. How long does it take the clerk to process 1 application? h

OpenStudy (anonymous):

@EclipsedStar

OpenStudy (anonymous):

@bohotness

Mehek (mehek14):

How many hours?

OpenStudy (anonymous):

3/4 hours

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@TeenWolfGirl

OpenStudy (anonymous):

Sorry idk

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

@lovinbabe We set up a proportion.

OpenStudy (igreen):

\(\dfrac{0.75}{9} = \dfrac{1}{x}\) Cross multiply: 0.7 * x = 9 * 1 Simplify and solve for 'x'